# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ 3.4 Group 4: Other reviews related to AI and technology in the health domain
A small number of reviews take a broader lens and consider different aspects of AI and technology in the health domain. This includes the concept of Healthcare 4.0, a term referring to the increasing digitisation of the healthcare industry. The reviews by Tortorella et al. [62] and Jayaraman et al. [63] broadly cover Healthcare 4.0, and highlight health monitoring systems that use IoT and sensors. However, only the review by Jayaram et al. [63] mentions ontologies and other knowledge representation techniques. A recent review by Kumar et al. [64] on AI in healthcare mentions IoT and knowledge graphs, but neither of these are the focus of the review. The review by Behera et al. [65] focuses on techniques used to create healthcare systems modeled on human cognitive processes such as perception and thought. They highlight cognitive IoT as a future research direction through wearable sensors, while also mentioning Semantic Web technologies for knowledge representation. However, neither the Semantic Web technologies nor sensors are discussed in detail.
@@ 3.5 Summary
Table 1 summarises the related reviews. The current study differs from existing work by focusing on the use of sensors and Semantic Web technologies for personal health monitoring, with both sensor data and Semantic Web technologies being primary points of focus. Additionally, the majority of the related reviews and surveys do not take a systems perspective, whereas this study highlights how the different system components are integrated and discusses the development methodologies and tools, evaluation approaches, and architectures of the included systems.
@@ 4.1 Objectives and reporting strategy
    To determine the role that Semantic Web technologies play in addressing these challenges.
@@
Given the goal and objectives of this work, a mapping study was the most appropriate approach. Although systematic mapping studies are similar to systematic literature reviews in terms of the systematic process of searching for and selecting studies, literature reviews aim at synthesizing evidence while mapping studies structure a research area through classification and categorisation in order to discover research trends [66]. The study was conducted and is reported using the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) [67] framework. To further ensure the quality of the study, we adhered to the following quality assessment criteria as described by Kitchenham et al. [68]:
@@
    “The authors have explicitly defined quality criteria and extracted them from each primary study”: The systems are analysed in Section 5 based on the seven identified challenges and the criteria is outlined in Table 8. Additionally, the quality of each system is assessed and discussed in Section 6 based on criteria outlined in Table 13.
@@
    “Information is presented about each paper so that the data summaries can clearly be traced to relevant papers”: A summary of all the included systems is shown in Table 4, with all systems fully cited. A GitHub repository [^12] https://github.com/mbithenzomo/semantic_phms_mapping_study has been created for this study, which includes copies of the relevant papers and other supplementary material.
@@ 4.2 Search strategy
Six digital libraries were searched between 9th and 12th February 2024: ACM Digital Library [^13] https://dl.acm.org, IEEE Xplore [^14] https://ieeexplore.ieee.org/Xplore, PubMed [^15] https://pubmed.ncbi.nlm.nih.gov, ScienceDirect [^16] https://www.sciencedirect.com, Scopus [^17] https://www.scopus.com, and Web of Science [^18] https://www.webofscience.com. For all libraries, abstracts, titles, and/or keywords were searched using terms related to the topic of the study, at the intersection of five areas: Semantic Web technologies, sensors, the health domain, monitoring, and systems. The search strings used are shown in Table 2. Boolean operators were used for a more specific search, although the ScienceDirect library had a limit on the number of Boolean operators that could be used per search. This library also did not allow the use of wildcard characters. Across all libraries, the results were filtered to only include literature published in or after 2012 to ensure a state of the art study. Additionally, where possible, the results were filtered to only include conference papers and journal articles published in English. This filtered out other types of literature such as surveys and reviews, books and book chapters, research abstracts, posters and conference proceedings, as well as articles written in languages other than English.
@@
Rayyan [69], an online tool for the management of systematic reviews, was used to facilitate the screening process. The initial digital library search yielded 960 records, which were then screened and assessed for eligibility. The resulting records, together with the related review articles discussed in Section 3, were then used to identify further potentially relevant studies. This was done using two online tools, Semantic Scholar [^19] https://www.semanticscholar.org and Connected Papers [^20] https://www.connectedpapers.com. The supplementary search results were filtered by publication year (in or after 2012), and the titles and abstracts were screened to identify records containing the relevant search terms. 62 additional records were identified through the supplementary search process, resulting in a total of 1,022 identified records.
@@ 4.3 Inclusion and exclusion criteria
This study includes only peer-reviewed journal articles and conferences papers in English. Only systems that incorporate semantic techniques were included for analysis, and those without a well-defined semantic technique as an integral component were excluded. Additionally, because a system consists of several integrated components, studies reporting the development of only one component (for example, an ontology) were excluded. Of particular interest are sensors that measure physiological data (that is, biosignals and vital signs) and/or physical activity data (for example, daily step count). Applications of sensors outside health monitoring , such as activity recognition, fitness, or nutrition, were excluded. Furthermore, systems that do not have an analysis, inferencing, or reasoning component were also excluded. These inclusion and exclusion criteria are summarised in Table 3.
@@
Table continued on next page.
Inclusion and exclusion criteria.
# Criteria Inclusion Criteria Exclusion Criteria
@@
C6 Semantic Web technologies The system incorporates Semantic Web technologies as an integral component. Semantic Web technologies are either poorly defined or do not form an integral component of the system.
C7 Health monitoring The system focuses on health monitoring. The system has a focus outside the health domain, or is related to health but does not focus on health monitoring (e.g. systems focusing solely on other areas such as activity recognition, sports, fitness, and nutrition).
@@ 4.4 Selection results
From the 1,022 identified records, 303 duplicates were removed resulting in 719 unique records. Next, preliminary screening was done by reviewing the title and abstract of each record. At this stage, records were excluded for reasons such as not being focused on the health domain or not involving health monitoring. We also found that a number of records had bypassed some of the filters that were applied in the initial identification stage, such as publication year and language. 533 records were excluded based on the title and abstract screening. The remaining 186 papers were read in full to determine if they still met the inclusion criteria. One reason for exclusion at this stage was if the system had been extended in later work and the extension was one of the systems being assessed. In such cases, the extension was included in the study while the previous work was excluded. Additionally, a small number of publications were excluded due to the full text being inaccessible without additional payment. Ultimately, 43 systems were selected for inclusion in this study. Figure 2 shows a PRISMA flow diagram illustrating the identification, screening, eligibility, and inclusion process.
Figure 2: PRISMA flow diagram outlining the selection process.
@@ 4.5 Summary of selected systems
A summary of the 43 selected systems is shown in Table 4, while Figure 3 shows the distribution of the systems according to the publication year. The year of publication ranges from 2012 to 2024, with 2021 being the most common. In terms of the application area, 25 focus on a particular disease or diseases, while the remaining 18 provide a solution for general health monitoring. Regarding the types of Semantic Web technologies used in the systems, nearly all of them make use of ontologies. The exceptions are the systems proposed by Yu et al. [70] and Zhou et al. [71] which use only knowledge graphs, and the one proposed by Xu et al. [72], which uses both linked data and a knowledge graph. Similarly, Reda et al. [73] use both linked data and an ontology, while Stavropoulos et al. [74] and Zafeiropoulos et al. [75] use both a knowledge graph and an ontology. Table 4 also provides an overview of other complimentary technologies and techniques used, as well as the architecture type of each system. These aspects are discussed in more detail in Sections 5 and  7 respectively.
With respect to the types of data used in the systems, all 43 systems collect physiological and other body data. 13 systems additionally incorporate environmental data from ambient sensors, while 19 consider data from existing health and medical records. Table 5 provides an overview of the types of sensor and non-sensor data used in the systems. A detailed analysis of the data sources, including sensor devices and existing datasets, is provided in Section 6.1.
@@
# System Year Application Semantic Technologies Other Technologies & Techniques Architecture Type
@@
1 Akhtar et al. [76] 2022 Parkinson’s Ontology Agents; CDL; Rules Layered; multi-agent
@@
2 Ali et al. [77] 2021 Diabetes; ABP Ontology ML; NLP; Rules Layered
3 Ali et al. [78] 2020 Heart disease Ontology ML; Rules Layered
4 Ali et al. [79] 2018 Diabetes Ontology FL; Queries; Rules Layered
@@
5 Alti et al. [80] 2022 Diabetes Ontology Agents; Queries; Rules Layered; multi-agent; service-oriented
6 Chatterjee et al. [81] 2021 Obesity Ontology Queries; Rules Modular
7 Chiang and Liang [82] 2015 General monitoring Ontology FL; Rules Modular
8 De Brouwer et al. [83] 2022 Headache disorders Ontology ML; Queries Modular
9 El-Sappagh et al. [84] 2019 Diabetes Ontology Queries; Rules Modular
10 Elhadj et al. [85] 2021 General monitoring Ontology Rules Layered
11 Esposito et al. [86] 2018 Arrhythmia Ontology FL; Rules Layered
12 Fenza et al. [87] 2012 General monitoring Ontology Agents; FL; Rules Layered; multi-agent; service-oriented
13 Garcia-Valverde et al. [88] 2014 General monitoring Ontology ML; Rules Unclear
14 Hadjadj and Halimi [89] 2021 General monitoring Ontology Queries; Rules Layered
15 Henaien et al. [90] 2020 General monitoring Ontology ML; Queries; Rules Layered
16 Hooda and Rani [91] 2020 Diabetes; heart disease Ontology Queries; Rules Modular
17 Hristoskova et al. [92] 2014 Heart failure Ontology Rules Service-oriented
18 Hussain and Park [93] 2021 Stroke Ontology ML; Queries; Rules Modular
[…]
21 Khozouie et al. [96] 2018 General monitoring Ontology Rules Modular
22 Kim et al. [97] 2014 General monitoring Ontology Queries; Rules Layered
23 Kordestani et al. [98] 2021 Kidney disease; skin disease Ontology ASP; Rules; BN Layered
24 Lopes de Souza et al. [99] 2023 Hypertension Ontology Queries; Rules Layered; modular
25 Mavropoulos et al. [100] 2021 General monitoring Ontology Agents; ML; NLP; Rules Layered; modular; single-agent
26 Mcheick et al. [101] 2016 Stroke Ontology BN Layered
27 Mezghani et al. [102] 2015 Diabetes Ontology ML; Queries; Rules Layered; service-oriented
28 Minutolo et al. [103] 2016 Arrhythmia Ontology FL; Rules Modular
29 Peral et al. [104] 2018 Diabetes Ontology ML; NLP; Rules Unclear
@@ 18 Hussain and Park [93] 2021 Stroke Ontology ML; Queries; Rules Modular
30 Reda et al. [73] 2022 General monitoring Linked Data; Ontology Queries; Rules Layered
31 Rhayem et al. [105] 2021 Gestational diabetes Ontology Queries; Rules Modular
32 Spoladore et al. [106] 2021 Diabetes; pulmonary disease Ontology Queries; Rules Layered
33 Stavropoulos et al. [74] 2021 Multiple sclerosis Knowledge graph; Ontology Rules Modular
34 Titi et al. [107] 2019 General monitoring Ontology Queries; Rules Layered
35 Vadillo et al. [108] 2013 General monitoring Ontology Agents Layered; multi-agent
36 Villarreal et al. [109] 2014 Diabetes Ontology None specified Layered
37 Xu et al. [72] 2017 General monitoring Linked data; Knowledge graph CBR; Queries Layered; service-oriented
38 Yu et al. [70] 2022 Paediatric asthma Knowledge graph ML; NLP; Rules Modular
39 Yu et al. [110] 2017 General monitoring Ontology Queries; Rules Layered
40 Zafeiropoulos et al. [75] 2024 Parkinson’s Knowledge graph; Ontology ML; Queries; Rules Modular
@@
41 Zeshan et al. [111] 2023 General monitoring Ontology Queries; Rules Modular
@@
42 Zhang et al. [112] 2014 General monitoring Ontology Rules Layered; modular
43 Zhou et al. [71] 2022 General monitoring Knowledge graph ML Modular

### v3

@@ 3.4 Group 4: Other reviews related to AI and technology in the health domain
A small number of reviews take a broader lens and consider different aspects of AI and technology in the health domain. This includes the concept of Healthcare 4.0, a term referring to the increasing digitisation of the healthcare industry. The reviews by Tortorella et al. [77] and Jayaraman et al. [78] broadly cover Healthcare 4.0, and highlight health monitoring systems that use IoT and sensors. However, only the review by Jayaram et al. [78] mentions ontologies and other knowledge representation techniques. More recent reviews, such as those by Rahman et al. [79] and Rashid and Nemati [80], have begun addressing the transition toward Healthcare 5.0, which emphasizes explainability and human-centricity in health systems. While both of these reviews mention the role of sensors, IoT, and AI in healthcare, neither mention the Semantic Web. Another review in this group is by Kumar et al. [81], who discuss AI in healthcare. Although they mention IoT and knowledge graphs, neither of these are the focus of the review. Lastly, the review by Behera et al. [82] focuses on techniques used to create healthcare systems modelled on human cognitive processes such as perception and thought. They highlight cognitive IoT as a future research direction through wearable sensors, while also mentioning Semantic Web technologies for knowledge representation. However, neither the Semantic Web technologies nor sensors are discussed in detail.
@@ 3.5 Summary
Table LABEL:related-reviews-table summarises the related reviews. The current study differs from existing work by focusing on the use of sensors and Semantic Web technologies for personal health monitoring, with both sensor data and Semantic Web technologies being primary points of focus. Additionally, the majority of the related reviews and surveys do not take a systems perspective, whereas this study highlights how the different system components are integrated and discusses the development methodologies and tools, evaluation approaches, and architectures of the included systems.
@@ 4.1 Objectives and reporting strategy
    To assess the role and limitations of Semantic Web technologies in addressing these challenges.
@@
    To conduct a comprehensive quality assessment of selected systems based on data and devices used, system and component development, evaluation rigour, and accessibility of research outputs.
  5. 5.
    To propose a reference architecture that provides guidance for the design and development of new systems.
  6. 6.
@@
The study was conducted and is reported using the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) [84] framework. To further ensure the quality of the study, we adhered to the following quality assessment criteria as described by Kitchenham et al. [85]:
@@
    “The authors have explicitly defined quality criteria and extracted them from each primary study”: The systems are analysed in Section 5 based on the seven identified challenges and the criteria is outlined in Table LABEL:challenges-aspects. Additionally, the quality of each system is assessed and discussed in Section 6 based on criteria outlined in Table LABEL:quality-critera.
@@
    “Information is presented about each paper so that the data summaries can clearly be traced to relevant papers”: A summary of all the included systems is shown in Table LABEL:selected-systems, with all systems fully cited. A GitHub repository [^14] https://github.com/mbithenzomo/semantic_phms_mapping_study has been created for this study, which includes copies and links of the selected papers and other supplementary material.
@@ 4.2 Search strategy
Six digital libraries were searched: ACM Digital Library [^15] https://dl.acm.org, IEEE Xplore [^16] https://ieeexplore.ieee.org/Xplore, PubMed [^17] https://pubmed.ncbi.nlm.nih.gov, ScienceDirect [^18] https://www.sciencedirect.com, Scopus [^19] https://www.scopus.com, and Web of Science [^20] https://www.webofscience.com. An initial search was conducted between 9th and 12th February 2024, and a second search was conducted between 14th and 15th August 2025 to ensure inclusion of the most recent literature. Abstracts, titles, and/or keywords were searched using terms related to the topic of the study, at the intersection of five areas: Semantic Web technologies, sensors, the health domain, monitoring, and systems. The search strings used are shown in Table 2. Boolean operators were used for a more specific search, although the ScienceDirect library had a limit on the number of Boolean operators that could be used per search. This library also did not allow the use of wildcard characters. To ensure a state of the art study, all results were filtered to only include literature published in or after 2012 during the first search, and between 2024 and 2025 during the second search, thus covering the period between 2012 and 2025. Additionally, where possible, the results were filtered to only include conference papers and journal articles published in English. This filtered out other types of literature such as surveys and reviews, books and book chapters, research abstracts, posters and conference proceedings, as well as articles written in languages other than English. The first search yielded 960 records, while the second search yielded a further 217 records, resulting in a total of 1,177 records from the digital library search.
@@
The selected results from the digital library search, together with the related review articles discussed in Section 3, were then used to identify further potentially relevant studies through a related paper search. This was done using two online tools, Connected Papers [^21] https://www.connectedpapers.com and Semantic Scholar [^22] https://www.semanticscholar.org. The first related paper search in February 2024 identified 62 records while the second one in August 2025 identified 22 records, resulting in 84 additional records. Thus, the entire search process yielded a total of 1,261 records, which were then assessed for eligibility. Rayyan [^23] https://www.rayyan.ai[86], an online tool for the management of systematic reviews, was used to facilitate the screening and assessment process. The search results from each phase of the search process are summarized in Table 3.
Table 3: Summary of search results from both the first (February 2024) and the second (August 2025) searches.
Search type Digital library/Online tool First search Second search Total
Digital library search ACM Digital Library 171 41 212
IEEE Xplore 202 29 231
PubMed 90 19 109
ScienceDirect 74 36 110
Scopus 202 35 237
Web of Science 221 57 278
Total 960 217 1,177
Related papers search Connected Papers 43 16 59
Semantic Scholar 19 6 25
Total 62 22 84
Total number of identified records 1,022 239 1,261
@@ 4.3 Inclusion and exclusion criteria
This study includes only peer-reviewed journal articles and conferences papers written in English. Further, we only include systems that incorporate one of the three Semantic Web technologies (i.e. an ontology, knowledge graph, or the explicit use of linked data) as an integral system component, with a clear description of the technical implementation. Systems lacking sufficient implementation details were excluded, as this would compromise the quality assessment. Additionally, because a system consists of several integrated components, studies reporting the development of only one component (for example, an ontology) were excluded. Of particular interest to this study are sensors that measure physiological data (that is, biosignals and vital signs) and/or physical activity data (for example, daily step count). Applications of sensors outside health monitoring , such as activity recognition, fitness, or nutrition, were excluded. Furthermore, systems that do not have an analysis, inferencing, or reasoning component were also excluded. These inclusion and exclusion criteria are summarised in Table LABEL:criteria.
@@
C6 Semantic Web technologies The system incorporates at least one of the three Semantic Web technologies as an integral component. Semantic Web technologies either do not form an integral component of the system or else their technical implementation is insufficiently described.
C7 Health monitoring The system focuses on personal health monitoring. The system has a focus outside the domain of human health, or is related to health but does not focus on health monitoring (e.g. systems focusing solely on other areas such as activity recognition, sports, fitness, and nutrition).
@@ 4.4 Selection results
From the 1,261 identified records, 366 duplicates were removed resulting in 895 unique records. Next, preliminary screening was done by reviewing the title and abstract of each record. At this stage, records were excluded for reasons such as not being focused on the health domain or not involving health monitoring. We also found that a number of records had bypassed some of the filters that were applied in the initial identification stage, such as publication year and language. 664 records were excluded based on the title and abstract screening. The remaining 231 papers were read in full to determine if they still met the inclusion criteria. One reason for exclusion at this stage was if the system had been extended in later work and the extension was one of the systems being assessed. In such cases, the extension was included in the study while the previous work was excluded. Additionally, a small number of publications were excluded for reasons such as the full text being inaccessible without additional payment or the article being retracted. Ultimately, 48 systems were selected for inclusion in this study. Figure 2 shows a PRISMA flow diagram illustrating the identification, screening, eligibility, and inclusion process. The diagram also provides details on the specific reasons for exclusion and the corresponding number of publications excluded at each stage.
Figure 2: PRISMA flow diagram outlining the selection process, combining the results from both the .
@@ 4.5 Summary of selected systems
A summary of the 48 selected systems is shown in Table LABEL:selected-systems, while Figure 3 shows the distribution of the systems according to the publication year. The year of publication ranges from 2012 to 2025, with 2021 being the most common. In terms of the application area, 29 focus on a particular disease or diseases, while the remaining 19 provide a solution for general health monitoring. Additionally, 17 of the systems mention elderly people as a target user [87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103]. Personal health monitoring systems can be classified into three development stages: research prototypes (functional systems validated on synthetic or existing datasets but not tested with real users), user validated (systems that have undergone user studies or real-world testing), and deployed (production-level systems that are currently operational). Of the 48 systems, a majority are research prototypes (36 systems), with 12 user validated systems and no deployed system.
Regarding the types of Semantic Web technologies used in the systems, nearly all make use of ontologies. The three exceptions are Yu et al. [104] and Zhou et al. [103], who use only knowledge graphs, and Xu et al. [105], who use linked data and knowledge graphs. Six systems use multiple Semantic Web technologies: Kilintzis et al. [97] and Reda et al. [106] combine linked data with ontologies; Stavropoulos et al. [99] and Zafeiropoulos et al. [107] combine knowledge graphs with ontologies; Xu et al. [105] combine linked data with knowledge graphs; and Ammar et al. [108] incorporate all three technologies. Table LABEL:selected-systems also provides an overview of other complimentary technologies and techniques used, as well as the architecture type of each system. These aspects are discussed in more detail in Sections 5 and  7 respectively.
@@
# System Year Application Development Stage Semantic Web Technologies Other Technologies & Techniques Architecture Type
@@
1 Akhtar et al. [109] 2022 Parkinson’s disease Prototype Ontology Agents; CDL; rules Layered; multi-agent
@@
2 Ali et al. [110] 2021 Diabetes; abnormal blood pressure Prototype Ontology ML; NLP; rules Layered
3 Ali et al. [111] 2020 Heart disease Prototype Ontology ML; rules Layered
4 Ali et al. [112] 2018 Diabetes User validated Ontology FL; queries; rules Layered
@@
5 Alti et al. [113] 2021 Diabetes Prototype Ontology Agents; queries; rules Layered; multi-agent; service-oriented
@@
6 Ammar et al. [108] 2021 Diabetes Prototype Linked data; KG; ontology Agents; queries; rules Layered; multi-agent
@@
7 Bampi et al. [87] 2025 General monitoring Prototype Ontology Rules; queries Modular
8 Chatterjee et al. [114] 2021 Obesity Prototype Ontology Queries; rules Modular
@@
9 Chiang and Liang [88] 2015 General monitoring Prototype Ontology FL; rules Modular
10 De Brouwer et al. [115] 2022 Headache User validated Ontology ML; queries Modular
11 El-Sappagh et al. [116] 2019 Diabetes Prototype Ontology Queries; rules Modular
12 Elhadj et al. [89] 2021 General monitoring Prototype Ontology Rules Layered
13 Esposito et al. [90] 2018 Arrhythmia User validated Ontology FL; rules Layered
14 Fenza et al. [117] 2012 General monitoring Prototype Ontology Agents; FL; rules Layered; multi-agent; service-oriented
15 Garcia-Moreno et al. [91] 2023 Frailty; dependence User validated Ontology ML Layered; service-oriented
16 Garcia-Valverde et al. [92] 2014 General monitoring Prototype Ontology ML; rules Unclear
17 Hadjadj and Halimi [118] 2021 General monitoring User validated Ontology Queries; rules Layered
18 Henaien et al. [93] 2020 General monitoring Prototype Ontology ML; queries; rules Layered
19 Hooda and Rani [94] 2020 Diabetes; heart disease Prototype Ontology Queries; rules Modular
20 Hristoskova et al. [119] 2014 Heart failure User validated Ontology Rules Service-oriented
21 Hussain and Park [120] 2021 Stroke User validated Ontology ML; queries; rules Modular
22 Ivașcu and Negru [95] 2021 General monitoring Prototype Ontology Agents; ML; queries; rules Modular; multi-agent
23 Ivașcu et al. [96] 2015 Mental & degenerative disorders Prototype Ontology Agents; rules Modular; multi-agent
24 Khozouie et al. [121] 2018 General monitoring Prototype Ontology Rules Modular
25 Kilintzis et al. [97] 2019 COPD User validated Linked data; ontology Rules; queries Layered; service-oriented
26 Kim et al. [122] 2014 General monitoring Prototype Ontology Queries; rules Layered
27 Kordestani et al. [123] 2021 Kidney disease; skin disease Prototype Ontology ASP; rules; BN Layered
28 Lopes de Souza et al. [124] 2023 Hypertension Prototype Ontology Queries; rules Layered; modular
29 Martella et al. [125] 2025 General monitoring Prototype Ontology Agents; rules; queries Layered; modular; multi-agent; service- oriented
30 Mavropoulos et al. [126] 2021 General monitoring User validated Ontology Agents; ML; NLP; rules Layered; modular; single-agent
31 Mcheick et al. [127] 2016 Stroke Prototype Ontology BN Layered
32 Mezghani et al. [128] 2015 Diabetes Prototype Ontology ML; queries; rules Layered; service-oriented
33 Minutolo et al. [129] 2016 Arrhythmia Prototype Ontology FL; rules Modular
34 Peral et al. [130] 2018 Diabetes Prototype Ontology ML; NLP; rules Unclear
35 Reda et al. [106] 2022 General monitoring Prototype Linked data; ontology Queries; rules Layered
36 Rhayem et al. [83] 2021 Gestational diabetes Prototype Ontology Queries; rules Modular
37 Spoladore et al. [98] 2021 Diabetes; COPD Prototype Ontology Queries; rules Layered
38 Stavropoulos et al. [99] 2021 Multiple sclerosis User validated KG; ontology Rules Modular
39 Titi et al. [100] 2019 General monitoring Prototype Ontology Queries; rules Layered
40 Vadillo et al. [101] 2013 General monitoring User validated Ontology Agents Layered; multi-agent
41 Villarreal et al. [102] 2014 Diabetes User validated Ontology None specified Layered
42 Xu et al. [105] 2017 General monitoring Prototype Linked data; KG CBR; queries Layered; service-oriented
43 Yu et al. [104] 2022 Paediatric asthma Prototype KG ML; NLP; rules Modular
44 Yu et al. [131] 2017 General monitoring Prototype Ontology Queries; rules Layered
45 Zafeiropoulos et al. [107] 2024 Parkinson’s disease Prototype KG; ontology ML; queries; rules Modular
46 Zeshan et al. [132] 2023 General monitoring User validated Ontology Queries; rules Modular
47 Zhang et al. [133] 2014 General monitoring Prototype Ontology Rules Layered; modular
48 Zhou et al. [103] 2022 General monitoring Prototype KG ML Modular